OpenStudy (anonymous):

i got three questions and that's it . simplification of the expression 4 over y exponent -3

OpenStudy (anonymous):

please help !

OpenStudy (anonymous):

\[\LARGE \left(\frac 4y\right)^{-3}\] this ? :)

OpenStudy (anonymous):

yes except the -3 is with the y ,

OpenStudy (anonymous):

its 4 over y ^-3

OpenStudy (anonymous):

Got it :) ...

OpenStudy (anonymous):

\[\LARGE \frac{4}{y^{-3}}=\cfrac{4}{\frac{1}{y^3}}=\cfrac{\cfrac41}{\cfrac{1}{y^3}}=\frac{4\cdot y^3}{1\cdot 1}=4y^3\]

OpenStudy (anonymous):

do you understand it ? :)

OpenStudy (anonymous):

next questions remember i need help with three questions Choose the correct simplification of the expression a^6b^3 over a^4b^2 i got choices A. A^2 / B B.A^10B^5 C.A^10/B^5 D.A^2B

OpenStudy (anonymous):

Find the attached image..You might understand it

OpenStudy (anonymous):

\[\LARGE \frac{a^6b^3}{a^4b^2}\] use this rule: \[\LARGE \frac{x^a}{x^b}=x^{a-b}\] can you do it ? :)

OpenStudy (anonymous):

let me try . :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what would be x ?

OpenStudy (anonymous):

\[\huge \frac{a^6b^3}{a^4b^2}=a^{6-4}b^{?-?}\] try this :)

OpenStudy (anonymous):

i got choice b

OpenStudy (anonymous):

No. that's wrong. Eliminate choice B , you have 3 choices remaining...

OpenStudy (anonymous):

okay . hmmm your a good teacher

OpenStudy (anonymous):

@Kreshnik, you're a very good tutor. I'm watching as @janayx3 learns.

OpenStudy (anonymous):

nooooooooooo it's D !

OpenStudy (anonymous):

ahha LOL :P @janayx3 which choice do you think is correct ? A C D ?? (check my last HINT ) I think you should know it !

OpenStudy (anonymous):

@janayx3 That's Correct. Well done :)

OpenStudy (anonymous):

because you have 6-4 , and that leaves 2 , then 3-2 =1 leave the variable b by its self

OpenStudy (anonymous):

... 100% true.. @Rudi19 Thank you. ;)

OpenStudy (anonymous):

@janayx3 Can you solve this? :) \[\LARGE \frac{x^5y^3}{xy^2}\] \[\LARGE A) x^4y^2\] \[\LARGE B) x^4y\] \[\LARGE C) x^3y^2\]

OpenStudy (anonymous):

It's Option B... I have to go. Good Night. :)

OpenStudy (anonymous):

goodnight :\